7. In two trials with identical starting concentrations of reactants, the rate increases by a factor of 3.27 when the temperature is increased from 400 K to 670 K. What is the activation energy for this reaction? Express your answer in kJ / mol.

8. The activation energy for a particular reaction is 4.8 kJ mol-1. The temperature in one trial reaction is 350 K. To what temperature must the reaction be heated to increase the rate constant by a factor of 2.5?
Answer #1

7)

ln (k2 / k1) = Ea / R [1 / T1 - 1/ T2]

ln (3.27) = Ea / 8.314 x 10^-3 [1/ 400 - 1/ 670]

Ea = 9.77

activation energy Ea = 9.77 kJ/mol

8)

ln (k2 / k1) = Ea / R [1 / T1 - 1/ T2]

ln (2.5) = 4.8 / 8.314 x 10^-3 [1/ 350 - 1/ T2]

T2 = 787 K

temperature = 787 K